A stock culture was diluted by adding 100ul into 900ul of media, then by adding 200ul into 800ul. 100ul of the final dilution was plated and 12 colonies were counted. what was the stock concentration?
Answer #1

Initial concentration = Number of colonies X Dilution factor

Dilution factor = 100/1000 X 200/1000

                     = 1/10 X 1/5 = 1/50

Plate dilution factor = 100/1000 = 1/10

Total dilution factor = 1/50 X 1/10 = 1/500

Initial concentration = 12 X 500

                              = 6000